OpenStudy (anonymous):

Write the following equation set equal to zero and in factored form. (x - 2)(x - 3) = 2

OpenStudy (anonymous):

x^2-5x+6=0 Is that it?

OpenStudy (anonymous):

these are the answer choises (x - 1)(x - 4) = 0 (x - 2)(x - 3) = 0 (x - 4)(x - 5) = 0

OpenStudy (anonymous):

@trumpetmaster7777

OpenStudy (anonymous):

oh like that. Hmm.

OpenStudy (anonymous):

yes. my mind is bogoling

OpenStudy (anonymous):

(x-2)(x-3)-2=0 Is the first step

OpenStudy (anonymous):

okay, do i combine like terms?

OpenStudy (anonymous):

Ya, lets try that and see what we come up with.

OpenStudy (anonymous):

Its the second one? right? (x - 2)(x - 3) = 0

OpenStudy (anonymous):

x^2-5x+4=0

OpenStudy (anonymous):

You need to foil out the first terms then you get x^2-5x+6=2 then subract the 2

OpenStudy (anonymous):

(x-1) (x-4)

OpenStudy (anonymous):

I gtg, hope you understand.

OpenStudy (anonymous):

Thank you very much, you got my mind work good this morning. And i do. Thanks again